I have been working on this problem forever and cannot figure it out. Can you help me and show me how you got the answer? John spent a 1/4 of his life growing up, 1/6 of his life in college, 1/2 of his life as a teacher. He spent his last 6 years in retirement. How old was he when he died? Thanks.
Answer by stanbon(75887)

John spent a 1/4 of his life growing up, 1/6 of his life in college, 1/2 of his life as a teacher. He spent his last 6 years in retirement. How old was he when he died?
----
Let "x" be his age at death.
----
Equation:
growing up: (x/4)
college:: (x/6)
teacher: (x/2)
retirement: 6 years
----
(x/4) + (x/6) + (x/2) + 6 = x
Multiply thru by 12 to get:
3x + 2x + 6x + 72 = 12x
x = 72 (his age at death)
============================
Cheers,
Stan H.
============